# TabArena: A Living Benchmark for Machine Learning on Tabular Data

### v2

@@ B.2 Included Datasets Details
Table B.1 presents a detailed overview for all datasets included in TabArena-v0.1. We further share all tasks and datasets as an OpenML suite (ID $457$, alias "tabarena-v0.1"). We share BibTex and LaTeX code to easily reference datasets from TabArena in our data curation repository: tabarena.ai/data-tabular-ml-iid-study.
Table B.1: Datasets included in TabArena-v0.1. ’Dataset (Task) ID’ represents the OpenML dataset and task IDs, ’name’ the dataset name, and Ref. the reference corresponding to the dataset. ’N’ represents the no. of samples, ’d’ the no. of features, ’C’ the no. of classes (- for regression tasks), and ’cat’ represents the percentage of features that are categorical. ’Subset’ indicates whether the dataset has been used for the sub-benchmarks focusing on TabPFNv2 (left) and TabICL (right).
@@ B.3 Noteworthy Observations from Curation
  • For various datasets, it was not possible to automate the selection process, because the metadata that would be required is not available. Therefore, given the current state of data repositories, we consider that automated curation procedures produce more biased results than careful manual curation. Finding out which splits are appropriate for a task, or whether the targets were created using deterministic functions, requires substantial effort and oftentimes, reading and understanding the papers where datasets were introduced. To still make the inclusion of datasets as objective as possible, we introduce a checklist for new datasets in Section D.3.
@@
  • The huge amount of tiny datasets is likely an artifact of a time when data collection was done at a much smaller scale than nowadays. Only four of the 142 tiny datasets for which we found a publication date were published later than 2010. Moreover, many of the tiny datasets seem to have originated in educational content, such as books or toy examples in tutorials.
@@
  • Of the 254 datasets with alternative versions listed in Figure 2, most are from the PMLB benchmark [44] and represent differently parameterized versions of artificially created datasets: 118 are Feynman equations and 62 are Friedman data generation functions.
@@ C.1 Implementation Framework Details
  • RandomForest and ExtraTrees use ordinal encoding for categorical variables. Missing values are imputed to 00.
@@
  • Linear uses one-hot-encoding, mean or median imputation (hyperparameter), standard scaling, and quartile transformation (hyperparameter).
  • KNN drops all categorical features and fills missing numerical values with 00. Moreover, it uses leave-one-out cross-validation instead of $8$-fold cross-validation. The leave-one-out cross-validation is natively implemented into the KNN model logic and allows for obtaining the validation predictions per sample very efficiently.
@@ C.2 Hyperparameter search spaces
  • The search spaces for CatBoost (Table C.1), LightGBM (Table C.2), XGBoost (Table C.3), RandomForest (Table C.4), and ExtraTrees (Table C.5) were determined based on experiments. We assessed a large set of hyperparameters inspired by the respective documentations as well as several papers [e.g., 37, 106, 20] and experimentally determined good ranges for them for tuning. We verified that the new search spaces outperform the original search spaces on TabRepo [37].
    For gradient-boosted trees, we use the implementation from AutoGluon with its early stopping logic and n_estimators=10_000. Compared to TabRepo, which used 300 overall estimators for Random Forest and ExtraTrees and used out-of-bag predictions for validation, we fit these models using 8-fold CV with 50 estimators per model, resulting in 400 estimators overall.
  • For EBM(Table C.6), we use a search space provided by the authors.
  • For RealMLP (Table C.7), we also use a search space provided by the authors. For the default parameters, we turn off label smoothing since we are not using accuracy as our evaluation metric, as recommended by [20].
  • For TabM (Table C.8) and ModernNCA (Table C.9), we use search spaces coordinated with the authors. For the batch size, we choose a training-set-size dependent logic following the TabM paper [9]:
[…]
  • FastaiMLP (Table C.10) and TorchMLP (Table C.11) were taken from AutoGluon, in dialogue with the maintainers/authors.
  • Linear (Table C.12) and KNN (Table C.13) were taken from TabRepo.
@@
  • For TabICL and TabDPT, we only use their default configurations. For TabICL and TabDPT, we use the newest checkpoint (see Section C.1), unlike the original paper.
@@
p Choice([2, 1])
C.3 TabArena Ensemble
The TabArena ensemble highlighted in Figure 6 was created by ensembling a portfolio, a set of hyperparameter configurations across models. Given a portfolio, we evaluate each of its models in sequence until a time limit is reached or all models have been evaluated. Then, we post-hoc ensemble [1] all evaluated hyperparameter configurations. For the sake of Figure 6, we simulated the TabArena ensemble using the result artifacts.
@@ C.3 TabArena Ensemble
Appendix D Using and Contributing to the Living Benchmark
D.1 Benchmarking with TabArena
@@
2from autogluon.core.models import AbstractModel
3from autogluon.features import LabelEncoderFeatureGenerator
4
5class CustomRandomForestModel(AbstractModel):
@@ 22 X=X
3 from tabrepo.utils.config_utils import ConfigGenerator
@@
2from tabrepo.benchmark.experiment import run_experiments_new
@@
9 model_experiments=model_experiments,
@@
2from tabrepo.paper.paper_runner_tabarena import PaperRunTabArena
@@ D.2 Contributing Models
    Public Model Implementation. The model must be implemented in the AbstractModel framework (see Section C.1), the code for this implementation must be publicly shared, and it must pass the default unit test for TabArena models. The implementation should represent a standalone model and not, for example, an ensembling pipeline of several existing models or sub-calls to other machine learning systems. We leave benchmarking for such pipelines, or in general, machine learning systems, to future iterations of TabArena.
@@
    Preprocessing and Hyperparameters. The implementation should specify model-specific preprocessing (see Section C.1). Moreover, the contributor must recommend default hyperparameters and a search space for hyperparameter optimization.
@@
    Model Verification. The maintainers of TabArena must have reviewed the source code of the model. In an ideal process, this review could also help the user to improve their model and implementation. In addition, the model should demonstrate promising results on TabArena-Lite. Moreover, if the contributor is not among the original authors of the model, the contributor (potentially in coordination with the maintainers of TabArena) shall reach out to the original authors to verify the implementation and its optimal intended usage. This may involve including the original author in GitHub issues, reviewing the pull request, or validating the results.
@@
D.3 Contributing Data: New Datasets and Curation Feedback
@@
Moreover, we also invite the community to add new datasets. For a new dataset to be added to TabArena, we outline the current template below:
@@
      1. If yes: Are tabular learning methods a reasonable solution compared to domain-specific methods? (If possible, provide a reference)
@@
The maintainers will verify the provided information and engage in review-like discussions if necessary. After verifying that the task is reasonable, the dataset will be included in the next benchmark version.
The above protocol outlines the user-driven process for adding new datasets. However, we welcome any suggestions for datasets that could be included in future versions of TabArena and where we, as maintainers, drive the process to add the dataset. For that, we welcome GitHub issues with the ‘Dataset Suggestion’ template, which includes: (1) a link to the raw data, and (2) the dataset license. The TabArena maintainers will review the suggested dataset by applying the outlined protocol and, if the criteria are met, include it in the next version of TabArena.
@@ D.3.1 Checklist Examples
D.4 Contributing Results: Leaderboard Submissions
@@
      Reproducible and documented code to obtain the results. We require users to start the process to add their new model to TabArena (as described in Section D.2) and to train and evaluate their approach using the provided TabArena benchmarking code.
@@
D.5 Running TabArena Models in Practice

### v3

@@ B.2 Included Datasets Details
Table˜B.1 presents the domain coverage of all datasets included in TabArena-v0.1.
Table˜B.2 presents a detailed overview for all datasets included in TabArena-v0.1. We further share all tasks and datasets as an OpenML suite [29, 61, 62] (ID $457$, alias "tabarena-v0.1"). We share BibTex and LaTeX code to easily reference datasets from TabArena in our data curation repository: tabarena.ai/data-tabular-ml-iid-study.
Table B.1: Domain Coverage of TabArena-v0.1. We present the domain name, along with the count and percentage of datasets from each domain, for all datasets in TabArena. We follow the domain names and categorization of the TALENT benchmark [36]. Note that TabArena does not have datasets for all categories defined by Ye et al. [36], as these categories conflict with our selection criteria.
Domain Count %
business & marketing 16 31.37%
finance 8 15.69%
chemistry & material science 6 11.76%
medical & healthcare 6 11.76%
biology & life sciences 5 9.80%
technology & internet 4 7.84%
physics & astronomy 3 5.88%
education 1 1.96%
environmental science & climate 1 1.96%
industry & manufacturing 1 1.96%
Table B.2: Datasets included in TabArena-v0.1. ’Dataset (Task) ID’ represents the OpenML dataset and task IDs, ’name’ the dataset name, and Ref. the reference corresponding to the dataset. ’N’ represents the no. of samples, ’d’ the no. of features, ’C’ the no. of classes (- for regression tasks), and ’% cat’ represents the percentage of features that are categorical. ’Subset’ indicates whether the dataset has been used for the sub-benchmarks focusing on TabPFNv2 (left) and TabICL (right).
@@ B.3 Noteworthy Observations from Curation
  • For various datasets, it was not possible to automate the selection process, because the metadata that would be required is not available. Therefore, given the current state of data repositories, we consider that automated curation procedures produce more biased results than careful manual curation. Finding out which splits are appropriate for a task, or whether the targets were created using deterministic functions, requires substantial effort and oftentimes, reading and understanding the papers where datasets were introduced. To still make the inclusion of datasets as objective as possible, we introduce a checklist for new datasets in Section˜E.3.
@@
  • The huge amount of tiny (fewer than 500 training samples) datasets is likely an artifact of a time when data collection was done at a much smaller scale than nowadays. Only four of the 142 tiny datasets for which we found a publication date were published later than 2010. Moreover, many of the tiny datasets seem to have originated in educational content, such as books or toy examples in tutorials.
@@
  • Of the 254 datasets with alternative versions listed in Figure˜2, most are from the PMLB benchmark [47] and represent differently parameterized versions of artificially created datasets: 118 are Feynman equations and 62 are Friedman data generation functions.
@@ C.1 Implementation Framework Details
  • RandomForest and ExtraTrees use ordinal encoding for categorical variables. Missing values are imputed to 0.
@@
  • Linear uses one-hot-encoding, mean or median imputation (hyperparameter), and standard scaling or quantile transformation (hyperparameter).
  • KNN uses different numerical and categorical feature encoding techniques as part of the search space. For numerical features, either z-standardization or quantile transformation is used. For categorical features, a ’cat_threshold’ parameter is defined determining whether categorical features are dropped, ordinal-encoded, or one-hot-encoded. If the no. of unique values is below the threshold, a feature is one-hot-encoded and otherwise ordinal-encoded. A value of 0 means that all categorical features are dropped. Missing numerical values are filled with 0. Moreover, KNN uses leave-one-out cross-validation instead of $8$-fold cross-validation. The leave-one-out cross-validation is natively implemented into the KNN model logic and allows for obtaining the validation predictions per sample very efficiently.
@@ C.2 Hyperparameter search spaces
  • The search spaces for CatBoost (Table˜C.1), LightGBM (Table˜C.2), XGBoost (Table˜C.3), RandomForest (Table˜C.4), and ExtraTrees (Table˜C.5) were determined based on experiments. We assessed a large set of hyperparameters inspired by the respective documentations as well as several papers [e.g., 37, 117, 20] and experimentally determined good ranges for them for tuning. We verified that the new search spaces outperform the original search spaces on TabRepo [37].
    For gradient-boosted trees, we use the implementation from AutoGluon with its early stopping logic and n_estimators=10_000. For Random Forest and ExtraTrees, TabRepo fit 300 trees on training+validation data and used out-of-bag predictions for validation. Since we want to allow bootstrap=False, which does not support out-of-bag predictions, we fit these models using 8-fold CV with 50 estimators per model, resulting in 400 estimators overall.
  • For EBM(Table˜C.6), we use a search space provided by the authors.
  • For RealMLP (Table˜C.7), we also use a search space provided by the authors. For the default parameters, we turn off label smoothing since we are not using accuracy as our evaluation metric, as recommended by [20].
  • For TabM (Table˜C.8) and ModernNCA (Table˜C.9), we use search spaces coordinated with the authors. For the batch size, we choose a training-set-size dependent logic following the TabM paper [9]:
    $\text{batch\_size }=\begin{cases}32&,N<2800\\ 64&,N\in[2800,4500)\\ 128&,N\in[4500,6400)\\ 256&,N\in[6400,32000)\\ 512&,N\in[32000,108000)\\ 1024&,N\geq 108000~.\end{cases}$
  • FastaiMLP (Table˜C.10) and TorchMLP (Table˜C.11) were taken from AutoGluon, in dialogue with the maintainers/authors.
  • Linear (Table˜C.12) and KNN (Table˜C.13) were taken from TabRepo. For KNN additional preprocessing was added to deal with varying numerical feature distributions and categorical features. For Linear, we used a log-uniform instead of uniform search space for the regularization parameter C.
@@
  • For TabICL and TabDPT, we only use their default configurations. For TabICL and TabDPT, we use the newest checkpoint (see Section˜C.1), unlike the original paper.
@@
p Choice([2, 1, 1.5])
scaler Choice(["standard", "quantile"])
cat_threshold Choice([0, 1, 5, 10, 20, 30, 50, 100, 1000000])
C.3 Cross-validation Ensembles
We employ $8$-fold cross-validation ensembles [27] within TabArena. Cross-validation ensembles function by ensembling models trained on different folds of the cross-validation process. In TabArena, we obtain $8$ fold models. We then ensemble the $8$ fold models by averaging their predictions, mirroring a bagging ensemble. The best configuration in the tuned regime is selected using the average out-of-fold performance over the $8$ folds.
Cross-validation ensembles are a powerful alternative to refitting to obtain a final model for deployment. The diversity from bagging the fold models can lead to better predictions. Furthermore, cross-validation ensembles are more efficient during training because there is no need to spend time refitting; however, they increase inference time.
C.4 Post-hoc Ensembling
Post-hoc ensembling (PHE) aims to combine a set of models previously evaluated on holdout data or by cross-validation during model selection (e.g., HPO) to improve performance over any single model [118, 43]. In particular, PHE relies on using data collected while evaluating models to build its ensemble, such as predictions on the validation data.
In practice, predictive machine learning systems most often [43] combine a set of models by aggregating their predictions with a weighted arithmetic mean whereby the weights of the models are commonly obtained using greedy ensemble selection (GES) [1, 119]. Likewise, multiple hyperparameter configurations of an individual tabular model can be ensembled with GES, as done in TabRepo [37] or by TabPFNv2 [5].
Post hoc ensembling with GES has four key advantages: 1) GES is very efficient due to reusing predictions on validation data previously collected while evaluating models [120, 43]; 2) GES optimizes a user-defined target metric using an anytime algorithm; 3) the final ensemble is usually small since GES produces sparse weight vectors [43, 44]; and 4) the predictive performance of post-hoc ensembling with GES is superior to the best individual model under mild assumptions [120, 121, 44, 43].
We build an ensemble of models using GES. In detail, we create an ensemble using the 200 hyperparameter configurations that were evaluated during the tuning process. To train the ensemble, we reuse the predictions on validation data that were computed during (inner) cross-validation. Then, we obtain a weight vector using GES [1, 119] for all configurations. Finally, we return the weighted average predictions of all non-zero-weighted configurations.
GES learns a weight vector $W=(w_{1},...,w_{m})$ to combine multiple models $f_{i}\in F$ from a pool of $m$-many models $F=(f_{1},...,f_{m})$ as $\sum_{i}w_{i}f_{i}$. GES ensures that $\forall i,0\leq w_{i}\leq 1$ and $\sum_{i}w_{i}=1$. The vector $W$ is learned via a greedy algorithm that runs for a fixed number of iterations NN ($N=40$ for TabArena). In each step $n\leq N$, GES finds $i$ such that increasing $w_{i}$ by $\frac{1-w_{i}}{n+1}$ and decreasing all other weights by $\frac{w_{j}}{n+1},\forall\,j\neq i$ most reduces the validation error.
C.5 TabArena Ensemble
The TabArena ensemble highlighted in Figure˜7 was created by ensembling a portfolio, a set of hyperparameter configurations across models. Given a portfolio, we evaluate each of its models in sequence until a time limit is reached or all models have been evaluated. Then, we post-hoc ensemble [1] all evaluated hyperparameter configurations. For the sake of Figure˜7, we simulated the TabArena ensemble using the result artifacts.
@@ C.5 TabArena Ensemble
Appendix D Evaluation Design Details
D.1 Elo Confidence Intervals
Suppose that the benchmark datasets $\mathcal{B}=(D_{1},\ldots,D_{51})$ are i.i.d. samples from an unknown dataset-generating distribution $P_{D}$. We want to compute a confidence interval for the “infinite-datasets Elo score” $\mathrm{Elo}(P_{D},A)$ of a method $A$, but can only compute finite-dataset Elo scores like $\operatorname{Elo}(\mathcal{B},A)$.
Similar to energies in physics, only relative differences between Elo scores are meaningful, as they predict win-rates between pairs of methods. We need to choose a reference point to obtain absolute Elo scores. We consider two variants:
  • $\operatorname{Elo}_{\mathrm{mean}}$: Center the mean Elo of all methods to 0.
  • $\operatorname{Elo}_{\mathrm{RF}}$: Center the Elo of default random forest to 1000.
For a ML method $A$, we compute 200 bootstrap subsamples $\tilde{\mathcal{B}}$ of $\mathcal{B}$. The 2.5% and 97.5% quantiles of $\operatorname{Elo}_{\mathrm{mean}}(\tilde{\mathcal{B}},A)$ yield an approximate 95% confidence interval $L_{\mathrm{mean}}(A),U_{\mathrm{mean}}(A)]$ for $\operatorname{Elo}_{\mathrm{mean}}(P_{D},A)$. Because of
$\operatorname{Elo}_{\mathrm{RF}}(\mathcal{B},A)=\operatorname{Elo}_{\mathrm{mean}}(\mathcal{B},A)+\Delta(\mathcal{B}),\quad\Delta(\mathcal{B}):=1000-\operatorname{Elo}_{\mathrm{mean}}(\mathcal{B},\text{RF (default)}),$
we report the shifted intervals $L_{A}+\Delta(\mathcal{B}),U_{A}+\Delta(\mathcal{B})]$, which are approximate confidence intervals for
$\operatorname{Elo}_{\mathrm{mean}}(P_{D},A)+\Delta(\mathcal{B})~.$
They are not good approximate confidence intervals for $\operatorname{Elo}_{\mathrm{RF}}(P_{D},A)$, because they do not factor in the randomness in the difference $\Delta(\mathcal{B})-\Delta(P_{D})=\operatorname{Elo}_{\mathrm{mean}}(P_{D},\text{RF (default)})-\operatorname{Elo}_{\mathrm{mean}}(\mathcal{B},\text{RF (default)})$. However, this term does not depend on $A$ and therefore does not affect the relative differences of results.
Discussion.
The use of $\operatorname{Elo}_{\mathrm{mean}}$-based confidence intervals explains why the confidence interval for RF(default) does not have length zero: We use a shift that shifts RF(default) to 1000 on this specific benchmark $\mathcal{B}$, but that would not shift it to 1000 on the ground-truth distribution $P_{D}$. We use $\operatorname{Elo}_{\mathrm{mean}}$ instead of $\operatorname{Elo}_{\mathrm{RF}}$ for confidence intervals because centering the mean instead of a weak method produces less variance and therefore smaller confidence intervals for strong methods. As a consequence, plots showing confidence intervals for absolute Elo values allow stronger conclusions about significance of relative differences in Elo values.
D.2 Sources of Randomness
The comparison and evaluation of models in TabArena is affected by various sources of randomness. The results in TabArena are affected by the following sources of randomness.
  • Model Randomness, resulting from: initialization, training, non-deterministic computations (on GPU or due to precision), inner validation splits (e.g., for early stopping), hyperparameter configurations, and the sampling of the hyperparameter configurations.
  • Data Randomness, resulting from: the selected datasets, the inherent sampling bias of each data set, and the partitions used for repeated outer cross-validation.
  • Evaluation Randomness, resulting from: metric calculation, and the precision of calculating metrics when the metric is used for ranking or normalization.
We guard against data randomness affecting our results by repeating our experiments several times per dataset. We partially guard against model randomness through repeating experiments and using many random configurations. Nevertheless, we do not fully guard against it, as we use a fixed random seed for models [^5] Future versions of TabArena will no longer use a fixed random seed. and a static set of random configurations. We guard against evaluation randomness by using 100 bootstrapping rounds and a stable Bradley-Terry Elo implementation.
D.3 Environmental Impact of TabArena
We are regrettably aware of the negative environmental impact of TabArena. We had several discussions about trading off environmental impact and research, but we have not converged on an official conclusion to this (philosophical) topic. To share some insights from our discussions, we provide several thoughts about the positive environmental impact of TabArena below. We argue that some of our key contributions will, over time, offset the negative environmental impact of running a large-scale benchmark.
  • We save and share the predictions (along with other metadata) of all models, allowing others to avoid wasting energy by rerunning our experiments to obtain the predictions for future studies.
  • We only simulate post-hoc ensembling on the saved predictions and thus save the computation overhead that could come from post-hoc ensembling, cf. [121, 43, 44].
  • We impose time limits on the training time of a model (per split). Thus, we avoid running configurations of models that would potentially take a very long time to converge, without improving predictive performance. A tighter time limit would save more energy.
  • The large (one-time) cost of running TabArena enables us to find more efficient and better models. Moreover, TabArena enables us to identify portfolios that improve the Pareto frontier in terms of both quality and efficiency. These portfolio configurations can be (and already are) implemented in predictive machine learning systems widely used in industry (e.g., AutoGluon), with applications that result in compute usage far exceeding that of the TabArena benchmark.
As a result of these contributions, the negative environmental impact of TabArena may be offset through its future applications, resulting in a net reduction of compute usage and a positive environmental impact.
Appendix E Using and Contributing to the Living Benchmark
E.1 Benchmarking with TabArena
@@
2import pandas as pd
3from autogluon.core.models import AbstractModel
4from autogluon.features import LabelEncoderFeatureGenerator
5
6class CustomRandomForestModel(AbstractModel):
@@ 23 X=X
3 from tabarena.utils.config_utils import ConfigGenerator
@@
2from tabarena.benchmark.experiment import run_experiments_new
@@
9 model_experiments=methods,
@@
2from tabarena.paper.paper_runner_tabarena import PaperRunTabArena
@@ E.2 Contributing Models
    Public Model Implementation. The model must be implemented in the AbstractModel framework (see Section˜C.1), the code for this implementation must be publicly shared, and it must pass the default unit test for TabArena models. The implementation should represent a standalone model and not, for example, an ensembling pipeline of several existing models or sub-calls to other machine learning systems. We leave benchmarking for such pipelines, or in general, machine learning systems, to future iterations of TabArena.
@@
    Preprocessing and Hyperparameters. The implementation should specify model-specific preprocessing (see Section˜C.1). Moreover, the contributor must recommend default hyperparameters and a search space for hyperparameter optimization.
@@
    Model Verification. The maintainers of TabArena must have reviewed the source code of the model. In an ideal process, this review could also help the user to improve their model and implementation. In addition, the model should (at least) demonstrate promising results on TabArena-Lite. Moreover, if the contributor is not among the original authors of the model, the contributor (potentially in coordination with the maintainers of TabArena) shall reach out to the original authors to verify the implementation and its optimal intended usage. This may involve including the original author in GitHub issues, reviewing the pull request, or validating the results.
@@
E.3 Contributing Data: New Datasets and Curation Feedback
@@
Moreover, we also invite the community to add new datasets and welcome any suggestions for datasets that could be included in future versions of TabArena. For a new dataset to be added to TabArena, there are two alternative processes: A maintainer-driven process and a user-driven process.
In the maintainer-driven process, we welcome GitHub issues with the ‘Dataset Suggestion’ template, which includes: (1) a link to the raw data, and (2) the dataset license. The TabArena maintainers will review the suggested dataset by applying the protocol outlined below and, if the criteria are met, include it in the next version of TabArena.
The user-driven process targets users with a high level of knowledge about the suggested dataset and requires users to follow our dataset inclusion template. We outline the current template below:
@@
      1. If yes: Are tabular learning methods a reasonable solution compared to domain-specific methods? (If possible, provide a reference).
@@
The maintainers will verify the provided information and engage in discussions if required. After verifying that the task is reasonable, the dataset will be included in the next benchmark version.
@@ E.3.1 Checklist Examples
E.4 Contributing Results: Leaderboard Submissions
@@
      Reproducible and documented code to obtain the results. We require users to start the process to add their new model to TabArena (as described in Section˜E.2) and to train and evaluate their approach using the provided TabArena benchmarking code.
@@
E.5 Running TabArena Models in Practice